OpenStudy (anonymous):

Solve using substitution:3x − y = 44x − 2y = 12 Answer (0, -4) (3, 5) (-2, -10) (6, 2)

OpenStudy (anonymous):

no

OpenStudy (ajprincess):

44x-6x+44=12 Can u find x solving the above equation @StudentAthlete?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

?

OpenStudy (ajprincess):

Really sorry. I made a mistake earlier. 3x-y=12 44x-2y=12 Solve ur first equation for y. Then u will get y=3x-12 Substitute this in ur second equation and u will get 44x-2(3x-12)=12 44x-2*3x-2*(-12)=12 44x-6x+24=12 Find x. Then substitute the value of x in y=3x-12 to find y.

OpenStudy (ajprincess):

R u sure u typed ur question correctly?

OpenStudy (anonymous):

your no help

OpenStudy (ajprincess):

none of your options seem to fit the question

OpenStudy (anonymous):

thats wat i copy n pasted

OpenStudy (ajprincess):

:3x − y =? Is there anythng after the equal sign?